OpenStudy (anonymous):

PLEASEE HEL ILL GIVEE MEDALSS ! Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

if two lines are perpndicular then product of there slopes is equal -1 so to find slope of desired line take slope of this line and then its -ve reciprocal..as here slope=2 so for desired line slope=-1/2 and then use this (y-y1)=m(x-x1) just plug in values and solve..